The brain volumes (cm3) of 20 brains have a mean of 1052.9 cm3 and a standard deviation of 121.3 cm3. Use the given standard deviation and the range rule of thumb to identify the limits separating values that are significantly low or significantly high. For such data, would a brain volume of 1245.5 cm3 be significantly high?


1
Expert's answer
2022-02-04T03:47:50-0500

Mean = 1161.2

Standard deviation = 122.7

The range rule of thumb indicates that the usual values are within 2 standard deviation from the mean.

Hence lower limit = Mean 2 x Standard deviation

=1161.2 - 2x122.7

=1161.2 - 245.4

=915.8

Upper Limit = Mean 2 x Standard deviation

=1161.2 + 2x122.7

=1161.2 + 245.4

=1406.6

The limit of values is from 915.8 to 1406.6


A value below 915.8 is considered low and a value above 1406.6 is considered high


1446.6 is more than the upper limit. Thus it is high for a brain.